OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles.A = 56°, a = 12, b = 14

OpenStudy (anonymous):

@VeritasVosLiberabit

OpenStudy (anonymous):

here are the options :)

OpenStudy (anonymous):

 Here are two triangles where a and b are placed in different ways. Two angles are shown for each where the triangle could be 56 degrees. Use trig functions to find what angles the triangle would be based on the different triangle configurations

OpenStudy (anonymous):

can you walk me through it?

OpenStudy (anonymous):

\[a=bcos(\beta)\] \[\cos ^{-1}(\frac{ a }{ b })=\beta \] \[\cos ^{-1}(\frac{ 12 }{ 14})\] \[180-(56+\cos ^{-1}(\frac{ 12 }{ 14 }))\]

OpenStudy (anonymous):

solve this to get the angles for one triangle

OpenStudy (anonymous):

124-cos^(-1)(6/7)?

OpenStudy (anonymous):

or about 123.5

OpenStudy (anonymous):

hold on I'm not getting the correct answer for these angles

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

These answers correspond to this problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I have not gotten a single same answer here. I'm not sure what's going on here.

OpenStudy (anonymous):

can you help me with a different problem then?

OpenStudy (anonymous):

sure... but this one is really strange or something is off

OpenStudy (anonymous):

Ya i couldn't figure it out either.

OpenStudy (anonymous):

That is really a strange one, you may want to check this one with your instructor